# High-Capacity and Low-PAPR BICM-OFDM Systems Using Non-Equiprobable and Non-Uniform Constellation Shaping With Clipping and Filtering

### v1

@@ II-A Geometric Design Based on Truncated Gaussian CDF
We first review an $M$-ary PAM constellation design based on the truncated Gaussian distribution [23]. Due to the symmetry of constellation, it is sufficient to consider the constellation points within the positive region. Let $\mathcal{M}_{+}\triangleq\{1,2,\cdots,M/2\}$ denote a set of constellation indices in this region. We define the set of corresponding constellation points as $\mathcal{A}_{+}\triangleq\{a_{k}\in{\mathbb{R}}_{+}\,|\,k\in{\mathcal{M}}_{+}\}$, with the elements ordered in the ascending order, i.e., $0<a_{1}\leq a_{2}\leq\cdots\leq a_{M/2}$. The constellation points are labeled by the binary reflected Gray code (BRGC), where the adjacent points differ only by a single bit. The most significant bit (MSB) is assigned as the sign bit such that $a_{k}$ and $a_{k}$ share the identical bit labels except for the MSB. Without loss of generality, we assume that MSB is 0 for the elements of $\mathcal{A}_{+}$. The resulting constellation including the negative region is given by $\mathcal{A}=\mathcal{A}_{+}\cup\mathcal{A}_{-}$, where $\mathcal{A}_{-}\triangleq\{-a_{k}\,|\,k\in{\mathcal{M}}_{+}\}$.
Since each PAM constellation point is selected with an equal probability, for a given target probability distribution, their locations are determined such that their cumulative distribution function (CDF) values are evenly distributed. Specifically, since the initial target distribution is Gaussian, we consider a random variable $X\sim{\cal N}\left(0,\frac{1}{2}\right)$. The CDF of $X$ is given by
@@
where $\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$ is the error function. Since the range of $X$ is unbounded, we truncate its range within [b,b][-b,b] for a given parameter $b>0$. Accordingly, the range of the CDF in the positive region should be modified as $\left[\frac{1}{2},F_{X}(b)\right]$. Thus, the tentative value $\hat{a}_{k}$, which corresponds to the point $a_{k}$ before energy normalization, is determined by the following equality:
@@
is the integer value unique to the $k$th constellation point.
Let us define $\rho$ as
$\rho\triangleq\operatorname{erf}(b),\quad\rho\in(0,1].$ (4)
We may then express $\hat{a}_{k}$ in (2) using (1) as
@@
where $\phi_{M}:\mathbb{N}\to\mathbb{R}_{+}$ is a constellation mapper for a given parameter $\rho$.
Finally, scaling the magnitude of $\hat{a}_{k}$ in accordance with the energy constraint leads to the desired constellation point $a_{k}$:
$a_{k}=\sqrt{\frac{E_{s}}{2}}\frac{\hat{a}_{k}}{\sqrt{\frac{2}{M}\sum_{k=1}^{M/2}\hat{a}^{2}_{k}}},$ (6)
where $E_{s}$ is the energy in terms of QAM symbol.
This design algorithm uniquely determines the resulting constellation using a single parameter $\rho\in(0,1]$ for a given modulation order $M$. When $\rho=1$ (i.e., without truncation), the resulting constellation coincides with the Gaussian-like PAM proposed in [21]. In contrast, we may express the Maclaurin series of $F_{X}(x)$ up to the first order as
@@
which indicates that $F_{X}(x)$ linearly increases (i.e., $X$ becomes a uniform distribution) for $X\in[-b,b]$ in the limit of $b\to 0$ ($\rho\to 0$). For this reason, we denote $\rho=0$ as the standard PAM in what follows.
Figure 1: Constellation point selection based on equiprobable partitioning of the Gaussian CDF ($\rho=1$) with $M=32$. Due to the symmetry, only the positive region is shown.
The relationship between the CDF values and the selected constellation points $\{\hat{a}_{k}\}$ in the case of $M=32$ and $\rho=1$ (i.e., without range truncation) is illustrated in Fig. 1. As shown in the figure, the values $\phi_{M}(2(k-1);\rho)$ and $\phi_{M}(2k;\rho)$ serve as the boundary of the $k$th equiprobable sub-region, which are depicted as the corresponding gray dashed lines. Each sub-region is assigned a single bit label, and the representative constellation point $\hat{a}_{k}$ is then selected to match the midpoint CDF value (i.e., $\phi_{M}(l_{k};\rho)$) of its corresponding region.
@@ II-B Constellation Point Merging Process
Since the optimal constellation cardinality that maximizes the BMI varies depending on the channel SNR, we consider dynamically adjusting the number of constellation points by incorporating a constellation point merging strategy into the truncated Gaussian-based design described above.
We define $N$ as the number of distinct points in $M$-ary PAM symbol with $N\leq M$, which means that a PAM constellation with $M$ distinct labels has only $N$ distinct constellation points. We denote this constellation as $(M,N)$-PAM in this work. The construction of $(M,N)$-PAM is given by the following recursive principle: adjacent point pairs in $\hat{\mathcal{A}}_{+}\triangleq\{\hat{a}_{k}\,|\,k\in{\mathcal{M}_{+}}\}$ with the minimum Euclidean distance are merged sequentially until the number of distinct points reaches the specified size $N$. The merging of the two symmetrical pairs in the positive and negative regions occurs simultaneously, leading to a reduction of two constellation points at each merging step. Therefore, for an initial modulation order of $M$, the procedure can be executed a maximum of M/21M/2-1 times in principle, ultimately leading to the standard BPSK constellation (when $N=2$).
Each merging operation combines the corresponding sub-regions into a new, unified sub-region, and the representative constellation point is placed to match the midpoint CDF value of this modified sub-region. Now, suppose that for a given parameter $\rho$, the constellation points $\hat{a}_{i}=\phi_{M}(l_{i};\rho)$ and $\hat{a}_{j}=\phi_{M}(l_{j};\rho)$ (with $i\neq j$) have the minimum Euclidean distance at a certain merging step. Then, these two points are merged into a single point by updating their indices and amplitude values as222 Since the initial indices $\{l_{k}\}$ are set to be positive odd numbers according to (3), the first merging step always applies to the two positive odd numbers. As a result, the merged index $l^{\prime}$ is guaranteed to be a natural number.
@@
After this process, the multiple bit labels originally associated with $\hat{a}_{i}$ and $\hat{a}_{j}$ are mapped to the identical signal amplitude. The proposed constellation shaping algorithm for a given pair of initial modulation order $M$ and the number of distinct points $N$ after the merging process is summarized in Algorithm 1, where $\mathcal{L}$ corresponds to the set of the constellation indices in the positive region that have unique constellation point locations. The algorithm thus terminates when $|\mathcal{L}|=N/2$.
Figure 2: Constellation point selection based on non-equiprobable partitioning of the Gaussian CDF ($\rho=1$) with $M=32$ and $N=10$.
@@ II-C Example
TABLE I: The $l_{k}$ indices of constellation points $\hat{\mathcal{A}}_{+}$ for the $(32,N)$-PAM with $\rho=1$ (no truncation). The second row indicates the bit labels. The initial amplitude values are shown in the third row, while the values after the merging process are shown in brackets. The number in the leftmost column represents that of distinct points $N$.
@@
30 2 [.055] 5 7 9 11 13 15 17 19 21 23 25 27 29 31
28 2 6 [.168] 9 11 13 15 17 19 21 23 25 27 29 31
26 2 6 10 [.284] 13 15 17 19 21 23 25 27 29 31
24 2 6 10 14 [.410] 17 19 21 23 25 27 29 31
22 2 6 10 14 18 [.549] 21 23 25 27 29 31
20 2 6 10 14 18 22 [.714] 25 27 29 31
18 4 [.111] 10 14 18 22 25 27 29 31
16 4 12 [.346] 18 22 25 27 29 31
14 4 12 18 22 26 [.932] 29 31
12 4 12 20 [.627] 26 29 31
10 8 [.225] 20 26 29 31
8 8 20 26 30 [1.317]
6 8 20 28 [1.085]
4 8 24 [.813]
2 16 [.477]
The integer representations $\{l_{k}\}$ of the constellation points for the positive region $\hat{\mathcal{A}}_{+}$ obtained by Algorithm 1 are listed in TABLE I in the case of $(32,N)$-PAM starting from $N=32$ to $2$, where the initial constellation points are derived from the original Gaussian distribution (i.e., $\rho=1$). The bit labels are based on BRGC. Therefore, the bit label of the negative region $a_{k}$ is identical to that of $a_{k}$ with MSB replaced by $1$. Focusing on the first merging step from $N=32$ to $30$, the initial constellation points near the origin (e.g., $\hat{a}_{1},\hat{a}_{2},\hat{a}_{3},\hat{a}_{4}$) exhibit very small Euclidean distances. When the adjacent point pair $\hat{a}_{1}$ and $\hat{a}_{2}$ is merged, they are modified into a single signal point but with two distinct bit labels. Then, these merged symbols become indistinguishable from each other, but they become more separable from the remaining symbols (e.g., $\hat{a}_{3},\hat{a}_{4}$). The conventional Gray labeling ensures that the interference from adjacent symbols results in only a single-bit error. However, in high-order constellations under low SNR conditions, the interference from non-adjacent symbols become likely as well, resulting in multiple-bit errors. The proposed constellation merging strategy effectively mitigates this issue by adaptively controlling the balance of density in terms of geometry and probability.
Fig. 2 illustrates the constellation point selection process based on the proposed merging scheme applied to the resulting constellation of Fig. 1, with the number of distinct points set to $N=10$. As shown in the figure, the original equiprobable sub-regions (gray dashed lines) are merged into larger sub-regions, depicted as blue dashed lines. Each representative constellation point is then selected in the same manner as that in Fig. 1, but based on the midpoint CDF value of the modified sub-region. This merging process corresponds to a non-equiprobable partitioning of the reference Gaussian distribution, according to the number of bit labels assigned to each constellation point.
We note that constellation point merging or many-to-one mapping in [26, 28, 29] refers to a low-complexity technique for implementing probabilistic shaping, where redundant bit labels are assigned to control the probability mass function (PMF) of the constellation to achieve shaping gain. In contrast, the proposed approach utilizes constellation point merging as an internal tool of geometric constellation design for the purpose of avoiding excessive density in the constellation. Therefore, this modification does not affect the complexity of the transmitter.
@@
Figure 3: Comparison between the proposed QAM consisting of two (32,20)-PAM constellations (shown as green circles) and the equivalent one-dimensional NUC adopted in ATSC 3.0 (shown as blue diamonds). Both constellations are optimized for BICM with an information rate of 8.333 bits per symbol (code rate of $3/5$).
The proposed QAM constellation points consisting of two $(32,20)$-PAM constellations where the parameter $\rho$ is optimized to maximize the BMI at the information rate corresponding to the code rate of $3/5$ (discussed in the next section) are depicted as green circles in Fig. 3. For comparison, those of the equivalent $1024$-point NUC adopted in ATSC 3.0 are plotted as blue diamonds. While the ATSC NUC individually optimizes all $32$ points per each dimension, many points around the origin are almost overlapping, resembling a smaller number of distinct points. Interestingly, the proposed constellation, systematically derived by Algorithm 1 and governed by only the two design parameters $N$ and $\rho$, exhibits a remarkably similar configuration to the NUC of ATSC 3.0. The performance analysis and optimization of these parameters leading to the above result are presented in the following section.
@@ III Performance Analysis
This section analyzes the performance that can be achieved by the proposed constellation design through parameter optimization, focusing on achievable gain and demapping complexity. We compare the results with other constellations found through numerical optimizations, which require significantly higher complexity than the proposed scheme. From Section II-B, for a given modulation order $M$, the proposed constellation is specified by the two parameters, $N$ and $\rho$. We denote the resulting constellation points located in both positive and negative regions by ${\cal A}\left(M,N,\rho\right)$ in what follows.
III-A Total BMI
To identify the constellations that are expected to perform well, it is necessary to evaluate their achievable information rates. In BICM systems, the achievable information rate is defined by the BMI [1]. For a given $M$-ary PAM constellation $\cal A$}, the total BMI at a specified channel SNR $\gamma$ is given by
$C\left(\mathcal{A},\gamma\right)=n-\sum_{i=0}^{n-1}{\mathbb{E}_{Y,b}\left[\log_{2}\frac{\sum_{x\in\mathcal{A}}p_{Y\mid X}(Y\,|\,x)}{\sum_{x\in\mathcal{A}_{b}^{(i)}}p_{Y\,|\,X}(Y\,|\,x)}\right]},$ (8)
measured in bits per dimension, where $n=\log_{2}M$ is the total number of binary channels and $\mathcal{A}_{b}^{(i)}\subset{\cal A}$ denotes the set of constellation points in the $i$th binary channel with the bit specified by $b\in\mathbb{F}_{2}$. The conditional probability density function (PDF) $p_{Y|X}(y\,|\,x)$ is defined as

### v2

@@ II-A Geometric Design Based on Truncated Gaussian CDF
We first review an $M$-ary PAM constellation design based on the truncated Gaussian distribution [kurihara2024design], where each PAM symbol has a binary label ${\bf b}\in{\mathbb{F}}_{2}^{m}$ with $m=\log_{2}M$. Due to the symmetry of the constellation, it is sufficient to consider the constellation points within the positive region. Let $\mathcal{M}_{+}\triangleq\{1,2,\cdots,M/2\}$ denote a set of constellation indices in this region. We define the set of corresponding constellation points as $\mathcal{A}_{+}\triangleq\{a_{k}\in{\mathbb{R}}_{+}\,|\,k\in{\mathcal{M}}_{+}\}$, with the elements ordered in the ascending order, i.e., $0<a_{1}\leq a_{2}\leq\cdots\leq a_{M/2}$. The constellation points are labeled by the binary reflected Gray code (BRGC) of $m$ bits, where the adjacent points differ only by a single bit. The most significant bit (MSB) is assigned as the sign bit such that $a_{k}$ and $a_{k}$ share the identical bit labels except for the MSB. Without loss of generality, we assume that MSB is 0 for the elements of $\mathcal{A}_{+}$. The resulting constellation, including the negative region, is given by $\mathcal{A}=\mathcal{A}_{+}\cup\mathcal{A}_{-}$, where $\mathcal{A}_{-}\triangleq\{-a_{k}\,|\,k\in{\mathcal{M}}_{+}\}$.
Since each PAM constellation point is selected with an equal probability, for a given target probability distribution, their locations are determined such that their CDF values are evenly distributed. Specifically, since the initial target distribution is Gaussian, we consider a random variable $X\sim{\cal N}\left(0,\frac{1}{2}\right)$. The CDF of $X$ is given by
@@
where $\operatorname{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$ is the error function. Since the range of $X$ is unbounded, we truncate its range within [b,b][-b,b] for a given parameter $b>0$. Accordingly, the range of the CDF in the positive amplitude region should be modified as $\left[\frac{1}{2},F_{X}(b)\right]$. Then, the tentative value $\hat{a}_{k}$, which corresponds to the point $a_{k}$ before energy normalization, is determined by the following equality:
@@
is the integer value unique to the $k$th constellation point. Let us define $\rho$ as
$\rho\triangleq\operatorname{erf}(b),\quad\rho\in(0,1]$ (4)
such that the CDF in the positive range of amplitude reduces to $\left[\frac{1}{2},\frac{1+\rho}{2}\right]$ according to (1). We may then express $\hat{a}_{k}$ in (2) using (1) as
@@
where $\phi_{M}:\mathbb{N}\to\mathbb{R}_{+}$ is a constellation mapper for a given parameter $\rho$. Finally, scaling the magnitude of $\hat{a}_{k}$ in accordance with the energy constraint leads to the desired constellation point $a_{k}$:
$a_{k}=\sqrt{\frac{E_{s}}{2}}\frac{\hat{a}_{k}}{\sqrt{\frac{2}{M}\sum_{\ell=1}^{M/2}\hat{a}^{2}_{\ell}}},$ (6)
where $E_{s}$ denotes the average transmit energy per complex-valued symbol. In our subsequent numerical calculation of the constellation points, we set $E_{s}=1$ for simplicity.
The above design algorithm uniquely determines the non-uniform PAM constellation using a single parameter $\rho\in(0,1]$ for a given modulation order $M$. When $\rho=1$ (i.e., without truncation), the resulting constellation coincides with the Gaussian-like PAM proposed in [boutros2018geometric]. In contrast, we may express the Maclaurin series of $F_{X}(x)$ up to the first order as
@@
which indicates that $F_{X}(x)$ linearly increases (i.e., $X$ becomes a uniform distribution) for $X\in[-b,b]$ in the limit of $b\to 0$ ($\rho\to 0$). For this reason, we denote $\rho=0$ as the standard uniform PAM in what follows.
(a) $\rho=1.0$
(b) $\rho=0.5$
Figure 1: Constellation point selection based on equiprobable partitioning of the Gaussian CDF with $M=32$. Due to the symmetry, only the positive region is shown. (a) is the case with $\rho=1.0$ (i.e., without truncation), while (b) corresponds to the case with $\rho=0.5$
The relationship between the CDF values and the selected constellation points $\{\hat{a}_{k}\}$ for the case of $M=32$ and $\rho=1$ (i.e., without truncation) is illustrated in Fig. 1(a). The values $\phi_{M}(2(k-1);\rho)$ and $\phi_{M}(2k;\rho)$ serve as the boundaries of the $k$th equiprobable subregion, which are depicted as the corresponding gray dashed lines. Each subregion is assigned a single binary label ${\bf b}\in\mathbb{F}_{2}^{m}$, and the representative constellation point $\hat{a}_{k}$ is then selected to match the midpoint CDF value (i.e., $\phi_{M}(l_{k};\rho)$) of the corresponding region. On the other hand, the case of $\rho=0.5$ is shown in Fig. 1(b), where the effective range of the original Gaussian CDF is now truncated to $\left[\frac{1}{2},\frac{1+\rho}{2}\right]=[0.5,0.75]$, as indicated by the red dashed line. Within this truncated range, we observe that the CDF increases almost linearly, and thus the resulting constellation points become more uniformly spaced, approaching those of the conventional uniform PAM constellation.
@@ II-B Constellation Point Merging Process
In coded modulation systems that do not utilize bit-wise interleaving, the achievable information rate, referred to as the constellation-constrained capacity, consistently increases with higher-order modulations. In BICM systems, the sum of the bit-wise mutual information (BMI) serves as the achievable information rate, and the constellation size that maximizes the total BMI depends on the channel SNR, indicating that high-order constellations may become suboptimal in low-SNR regimes [caire1998bit, zollner2013optimization]. Therefore, we next consider dynamically adjusting the number of constellation points by incorporating a strategy of merging constellation points into the truncated Gaussian-based geometrical design described above.
We define $N$ as the number of distinct points in an $M$-ary PAM symbol with $N\leq M$, which indicates that a PAM constellation with $M$ distinct labels has only $N$ distinct constellation points. We denote this constellation as $(M,N)$-PAM in this work. The construction of $(M,N)$-PAM is given by the following recursive principle: adjacent point pairs in $\hat{\mathcal{A}}_{+}\triangleq\{\hat{a}_{k}\,|\,k\in{\mathcal{M}_{+}}\}$ with the minimum Euclidean distance are merged sequentially until the number of distinct points reaches the specified size $N$. The merging of the two symmetrical pairs in the positive and negative regions occurs simultaneously, leading to a reduction of two constellation points at each merging step. Therefore, for an initial modulation order of $M$, the procedure can be executed a maximum of M/21M/2-1 times in principle, ultimately leading to the standard BPSK constellation (when $N=2$).
Each merging operation combines the corresponding subregions into a new, unified subregion, and the representative constellation point is placed to match the midpoint CDF value of this modified subregion. Now, suppose that for a given parameter $\rho$, the constellation points $\hat{a}_{i}=\phi_{M}(l_{i};\rho)$ and $\hat{a}_{j}=\phi_{M}(l_{j};\rho)$ (with $i\neq j$) have the minimum Euclidean distance at a certain merging step. Then, these two points are merged into a single point by updating their indices and amplitude values as [^3] Since the initial indices $\{l_{k}\}$ are set to be positive odd numbers according to (3), the first merging step always applies to the two positive odd numbers. As a result, the merged index $l^{\prime}$ is guaranteed to be a natural number.
@@
After this process, the multiple binary labels originally associated with $\hat{a}_{i}$ and $\hat{a}_{j}$ are mapped to the identical signal amplitude. The proposed shaping algorithm, which is summarized in Algorithm 1, applies to a specific pair of initial modulation order $M$ and the number of distinct points $N$ after the merging process, with $\mathcal{L}$ representing the set of the constellation indices in the positive amplitude region that have unique constellation point locations. The algorithm thus terminates when $|\mathcal{L}|=N/2$.
In what follows, a set of the resulting distinct PAM constellation points located in both positive and negative regions, after normalization through (6), will be denoted by ${\cal A}\left(M,N,\rho\right)$, which is uniquely determined by the two parameters, $N$ and $\rho$, for a given modulation order $M$. Furthermore, since each binary label is converted to a corresponding PAM symbol, we define the symbol mapping function as
$\varphi:{\mathbb{F}}_{2}^{m}\to{\cal A}\left(M,N,\rho\right)$ (8)
such that $\varphi(\bf b)$ represents the PAM symbol that has the binary label of length $m$ specified by $\bf b$}. We note that if $N<M$ the mapping function $\varphi$ will be characterized by many-to-one mapping.
Figure 2: Constellation point selection based on non-equiprobable partitioning of the Gaussian CDF ($\rho=1.0$) with $M=32$ and $N=10$.
@@ II-C Example
TABLE I: The $l_{k}$ indices of constellation points $\hat{\mathcal{A}}_{+}$ for the $(32,N)$-PAM with $\rho=1$ (no truncation). The second row indicates the binary labels. The initial amplitude values are shown in the third row, while the values after the merging process are shown in brackets. The number in the leftmost column represents NN (distinct points).
@@
[…]
The integer representations $\{l_{k}\}$ of the constellation points for the positive region $\hat{\mathcal{A}}_{+}$ generated by Algorithm 1 are listed in Table I in the case of $(32,N)$-PAM starting from $N=32$ to $2$, where the initial constellation points are derived from the original Gaussian distribution (i.e., $\rho=1$). Since the binary labels are based on BRGC, the bit labels of the negative region $a_{k}$ are identical to that of $a_{k}$ with the MSB replaced by $1$. Focusing on the first merging step from $N=32$ to $30$, the initial constellation points near the origin (e.g., $\hat{a}_{1},\hat{a}_{2},\hat{a}_{3},\hat{a}_{4}$) exhibit very small Euclidean distances. When the adjacent point pair $\hat{a}_{1}$ and $\hat{a}_{2}$ is merged, they are modified into a single constellation point that is assigned two distinct binary labels. Then, these merged symbols become indistinguishable from each other, while they become more separable from the remaining symbols (e.g., $\hat{a}_{3},\hat{a}_{4}$). As a result, the constellation behaves similarly to a lower-order constellation with fewer but more reliable symbols.
Fig. 2 illustrates the constellation point selection process based on the proposed merging scheme applied to the resulting constellation of Fig. 1(a), with the number of distinct points set to $N=10$. As shown in the figure, the original equiprobable subregions (gray dashed lines) are merged into larger subregions, depicted as blue dashed lines. Each representative constellation point is then selected in the same manner as that in Fig. 1, but based on the midpoint CDF value of the modified subregion. This merging process corresponds to a non-equiprobable partitioning of the reference Gaussian distribution, according to the number of binary labels assigned to each constellation point.
Considering the fact that the resulting constellation exhibits a non-uniform probability mass function (PMF), the proposed design can be viewed as a hybrid of geometric and probabilistic shaping. However, the non-uniformity in symbol probabilities arises from the rearrangement (i.e., geometrical design) of the Gray-labeled $M$ symbols, where multiple symbols are merged to have an identical amplitude. Thus, unlike conventional probabilistic shaping schemes that require a distribution matcher, the proposed shaping scheme imposes only the update of the bit-to-symbol mapping table in a PAM-based BICM system with the same modulation order $M$.
@@
Figure 3: Comparison between the proposed QAM consisting of two (32,20)-PAM constellations (shown as green circles) and the equivalent one-dimensional NUC adopted in ATSC 3.0 (shown as blue diamonds). Both constellations are optimized for BICM with an information rate of 6.0 bits per symbol (code rate of $3/5$). The size of the circles reflects their probability of occurrence.
The proposed NENU constellation consisting of two $(32,20)$-PAM constellations is depicted using green circles in Fig. 3, where symbols with higher probabilities are plotted with larger marks. The parameter $\rho$ is optimized based on the BMI for a target spectral efficiency of $6.0$ bits per symbol, corresponding to a code rate of $3/5$. For comparison, the equivalent $1024$-ary one-dimensional NUC adopted in ATSC 3.0 is depicted with blue diamonds. While the ATSC NUC individually optimizes all PAM amplitudes, many points near the origin are almost overlapping, resembling a smaller number of distinct points. The proposed constellation, which is systematically derived by Algorithm 1 and governed by only the two design parameters $N$ and $\rho$, exhibits a remarkably similar configuration to the NUC of ATSC 3.0. The performance analysis and optimization of these parameters leading to the above result are described in the following section.
@@ III Performance Analysis
This section analyzes the performance that can be achieved by the proposed constellation design through parameter optimization, focusing on achievable gain and demapping complexity. We compare the results with other constellations found through numerical optimizations, which require significantly higher complexity than the proposed scheme.
III-A Bit-wise Mutual Information
Let ${\mathcal{B}}_{b}^{(i)}\subset\mathbb{F}_{2}^{m}$ denote a set of binary labels with its $i$th bit specified by $b\in\mathbb{F}_{2}$. Let $\cal A$} denote a set of $M$-ary PAM constellation points that are mapped from the binary label through (8). The BMI of the $i$th binary channel at a specified channel SNR $\gamma$ over an AWGN channel is then expressed as [szczecinski2015bit, caire1998bit]
$C_{i}\left(\mathcal{A},\gamma\right)=1-\sum_{b\in\mathbb{F}_{2}}\sum_{{\bf b}\in\mathcal{B}_{b}^{(i)}}P({\bf b})\,$
$\times\int_{-\infty}^{\infty}p_{Y\mid X}(y\mid\varphi({\bf b}))\log_{2}\frac{\sum_{{\bf b}^{\prime}\in\mathbb{F}_{2}^{m}}p_{Y\mid X}(y\,|\,\varphi({\bf b}^{\prime}))}{\sum_{{\bf b}^{\prime\prime}\in\mathcal{B}_{b}^{(i)}}p_{Y\mid X}(y\,|\,\varphi({\bf b}^{\prime\prime}))}dy,$ (9)
where $P({\bf b})=\frac{1}{M}$ (i.e., the input binary labels are equiprobable) and the conditional probability density function (PDF) $p_{Y|X}(y\,|\,x)$ is defined as